# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v3

@@ 4.2.3. Workflow architecture
Communicative architectures can facilitate debate and conversation among workers (Du et al., 2023; Teevan et al., 2016; Wu et al., 2023a; Salisbury et al., 2017; Drapeau et al., 2016; Chen et al., 2019; Chang et al., 2017; Kobayashi et al., 2018; Liang et al., 2023) and allow workers to return feedback to the user (Kulkarni et al., 2012; Lasecki et al., 2015; Nebeling et al., 2016). Some workflows create roles for different workers, often based on roles held in real-life organizations (e.g., journalists, software development teams, etc.) (Valentine et al., 2017; Agapie et al., 2015; Retelny et al., 2014; Park et al., 2022). Some papers integrate these roles into a hierarchy of communication, with managers, midlevel managers, and different forms of communication between different levels of the hierarchy (Agapie et al., 2015; Retelny et al., 2014). Communication may be restrained through subtasks (Chang et al., 2017; Chen et al., 2019), free-form though email (Agapie et al., 2015; Valentine et al., 2017), or enabled through a shared interface (Lasecki et al., 2015; Lee et al., 2017; Nebeling et al., 2016; Kulkarni et al., 2014; Mahyar et al., 2018; Lasecki et al., 2013). Communication among workers may be synchronous or asynchronous. Synchronous workflows help enable collaboration (André et al., 2014; Chilton et al., 2014; Salisbury et al., 2017; Chilton et al., 2019) and can reduce latency (Bernstein et al., 2011). With asynchronous workflows or workflows with limited communication architectures, a working memory or state management can help workers access progress reports, what other workers think are good next steps, and the overarching goals (Valentine et al., 2017; Hahn et al., 2016; Lasecki et al., 2013; LaToza et al., 2014; Zhang et al., 2012; Hahn et al., 2016; Yang et al., 2022; Nair et al., 2023b; Wu et al., 2023a).
@@ 5. Strategies integrating the design space
Choosing tactic designs to support objectives is a non-trivial task, as combining different tactic types and implementations supports a huge range of possible workflows. However, through the thematic coding process we surfaced strategies by which tactics are used to support objectives (Figure 4). Specifically, these strategies (5.1) encourage diverse responses, (5.2) validate outputs from subtasks, (5.3) enable user guidance, (5.4) utilize adaptable architectures, (5.5) improve worker recruitment, (5.6) align a subtask’s required skills with a worker’s capabilities, (5.7) promote global context, and (5.8) ensure outputs meet a predetermined quality thresholds.
@@ 5.1. Diverse responses
To improve accuracy or creativity, source multiple responses to the same subtask with parallel, iterative, or debate based workflows (Figure 5).
@@
Motivation. Individual workers can give low quality outputs, affecting the quality of the overall output. Designers may want to improve robustness to low quality outputs, outliers, and spammers to support the accuracy of outcomes (Liem et al., 2011; Du et al., 2023; Zaidan and Callison-Burch, 2011; Huang et al., 2018; Lasecki et al., 2013; Huffaker et al., 2020; Reppert et al., 2023; Ambati et al., 2012; Hahn et al., 2016). For creative tasks, they may want to encourage divergent and diverse thinking (Liang et al., 2023; Wu et al., 2023a; Luther et al., 2015; Lasecki et al., 2015; Huang et al., 2017; Nebeling et al., 2016; Teevan et al., 2016).
@@
Implementation. Getting responses from multiple workers on a subtask can help support these goals. Crowdsourcing workflows and LLM chaining source diverse responses differently. Crowdsourcing does so by querying different people. Chaining leverages the stochastic nature of LLMs (e.g., with a higher temperature parameter setting) or or uses different models or prompt variants (including role-based prompts) to source a variety of responses (Park et al., 2022; Wu et al., 2023c; Arora et al., 2022). In some cases one can automatically calculate the optimal mix of LLMs for a task (Madaan et al., 2023a).
@@
Redundant architectures are the most direct way of sourcing diverse responses, but communicative architectures, such as those that structure debates, can source diverse responses as well. Some studies in crowdsourcing systematically study the impact of different architectures (e.g., parallel vs iterative) and architecture parameters (e.g., the number of iterations) on the diversity and quality of the outcomes (Goto et al., 2016; Little et al., 2010a; Kittur et al., 2011; André et al., 2014). However, it is unknown if these findings transfer to similar tactic choices with LLMs.
@@
Secondary benefits. There are some benefits of diverse responses that are solely talked about in the crowdsourcing literature, as they are dependent on having different people. Crowdsourcing brings up the benefit of having different viewpoints (Chen et al., 2019; Ambati et al., 2012; Lasecki et al., 2013), filling in each other’s deficits to create a more rounded final product (Kim and Monroy-Hernandez, 2016; Zhang et al., 2012), and sourcing perspectives from expert crowdworkers (Chilton et al., 2014) or from workers with various skill levels (Mahyar et al., 2018).
@@
Consequences. Sourcing diverse responses is inherently more costly because multiple workers complete the same subtask. Merging multiple responses can induce outcome quality errors: coherency for parallel architectures (Wu et al., 2022a) and cascading errors for iterative architectures (Goto et al., 2016).
@@
Examples. Examples of the diverse responses strategy include: Apparition (Lasecki et al., 2015), chains for multi-agent debate (Liang et al., 2023), and this exploration of ”human computation processes” (Little et al., 2010a).
@@
Title. Diverse responses, to improve accuracy or creativity, source multiple responses to the same subtask with parallel, iterative, or debate based workflows. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Robust to low quality responses, Coherency and cascading errors), Constraints (increase expense and latency), Accuracy (Robust to inaccurate outputs), Creativity (Improve with divergent thinking). Section 2: tactics used. Workers (LLMs, crowdworkers), Subtasks (all), Architectures (redundant, communicative).

### v4

@@ 4.2.3. Workflow architecture
Communicative architectures can facilitate debate and conversation among actors [C(Teevan et al., 2016; Drapeau et al., 2016; Chang et al., 2017; Kobayashi et al., 2018; Salisbury et al., 2017; Chen et al., 2019) L(Wu et al., 2023a; Du et al., 2023; Liang et al., 2023)] and allow workers to return feedback to the user [C(Nebeling et al., 2016; Kulkarni et al., 2012; Lasecki et al., 2015)]. Some workflows create roles for different actors, often based on roles held in real-life organizations (e.g., journalism teams, software development teams, etc.) [C(Agapie et al., 2015; Valentine et al., 2017; Retelny et al., 2014) L(Park et al., 2022)]. Some papers integrate these roles into a hierarchy of communication, with managers, midlevel managers, and different forms of communication between different levels of the hierarchy [C(Agapie et al., 2015; Retelny et al., 2014)]. Communication may be restrained through subtasks [C(Chang et al., 2017; Chen et al., 2019)], free-form though email [C(Agapie et al., 2015; Valentine et al., 2017)], or enabled through a shared interface [C(Kulkarni et al., 2014; Lasecki et al., 2015, 2013; Nebeling et al., 2016; Mahyar et al., 2018; Lee et al., 2017)]. Communication among actors may be synchronous or asynchronous. Synchronous workflows help enable collaboration [C(André et al., 2014; Chilton et al., 2019; Salisbury et al., 2017; Chilton et al., 2014)] and can reduce latency [C(Bernstein et al., 2011)]. In asynchronous workflows or workflows with limited communication architectures, a working memory can help actors access progress reports, what other actors think are good next steps, and the overarching goals [C(Valentine et al., 2017; Lasecki et al., 2013; LaToza et al., 2014; Hahn et al., 2016; Zhang et al., 2012) L(Wu et al., 2023a; Yang et al., 2022; Nair et al., 2023b)].
Reference |
Title |
Venue |
Year |
Quality Increase |
Quality Decrease |
Cost Increase |
Cost Decrease |
Objectivity |
Subjectivity |
Task General |
Task Specific |
Sequential |
Branching |
Redundant |
Dynamic |
Communicative |
Generate |
Evaluate |
Improve |
Focus |
Partition |
Merge |
Crowdworkers |
LLMs |
User |
| (Bernstein et al., 2010) | Soylent: a word processor with… | UIST | 2010 | ||||||||||||||||||||||
| (Dai et al., 2010) | Decision-Theoretic Control of… | AAAI | 2010 | ||||||||||||||||||||||
| (Little et al., 2010a) | Exploring Iterative and Parallel… | CKDD-HCOMP | 2010 | ||||||||||||||||||||||
| (Little et al., 2010b) | TurKit: human computation… | UIST | 2010 | ||||||||||||||||||||||
| (Kittur et al., 2011) | CrowdForge: crowdsourcing complex… | UIST | 2011 | ||||||||||||||||||||||
| (Ahmad et al., 2011) | The Jabberwocky Programming… | UIST | 2011 | ||||||||||||||||||||||
| (Marcus et al., 2011b) | Demonstration of Qurk: A Query… | SIGMOD | 2011 | ||||||||||||||||||||||
| (Marcus et al., 2011a) | Human-powered Sorts and… | VLDB | 2011 | ||||||||||||||||||||||
| (Bernstein et al., 2011) | Crowds in Two Seconds: Enabling… | UIST | 2011 | ||||||||||||||||||||||
| (Liem et al., 2011) | An Iterative Dual Pathway… | AAAI | 2011 | ||||||||||||||||||||||
| (Zaidan and Callison-Burch, 2011) | Crowdsourcing Translation:… | ACL | 2011 | ||||||||||||||||||||||
| (Ambati et al., 2012) | Collaborative Workflow for… | CSCW | 2012 | ||||||||||||||||||||||
| (Kulkarni et al., 2012) | Collaboratively crowdsourcing… | CSCW | 2012 | ||||||||||||||||||||||
| (Kittur et al., 2012) | CrowdWeaver: Visually Managing… | CSCW | 2012 | ||||||||||||||||||||||
| (Willett et al., 2012) | Strategies for Crowdsourcing… | CHI | 2012 | ||||||||||||||||||||||
| (Zhang et al., 2012) | Human computation tasks with… | CHI | 2012 | ||||||||||||||||||||||
| (André et al., 2013) | Community Clustering: Leveraging… | HCOMP | 2013 | ||||||||||||||||||||||
| (Lasecki et al., 2013) | Chorus: a crowd-powered… | UIST | 2013 | ||||||||||||||||||||||
| (Pietrowicz et al., 2013) | CrowdBand: An Automated… | HCOMP | 2013 | ||||||||||||||||||||||
| (Bragg et al., 2013) | Crowdsourcing Multi-Label… | HCOMP | 2013 | ||||||||||||||||||||||
| (Chilton et al., 2013) | Cascade: crowdsourcing taxonomy… | CHI | 2013 | ||||||||||||||||||||||
| (Dai et al., 2013) | POMDP-based control of workflows… | Elsiver | 2013 | ||||||||||||||||||||||
| (Kim et al., 2014) | Crowdsourcing step-by-step… | CHI | 2014 | ||||||||||||||||||||||
| (Lin et al., 2014a) | Crowdsourced Explanations for… | HCOMP | 2014 | ||||||||||||||||||||||
| (Pavel et al., 2014) | Video Digests: A Browsable,… | UIST | 2014 | ||||||||||||||||||||||
| (Retelny et al., 2014) | Expert crowdsourcing with flash… | UIST | 2014 | ||||||||||||||||||||||
| (Christoforaki and Ipeirotis, 2014) | STEP: A Scalable Testing and… | HCOMP | 2014 | ||||||||||||||||||||||
| (André et al., 2014) | Effects of Simultaneous and… | CHI | 2014 | ||||||||||||||||||||||
| (Kulkarni et al., 2014) | Wish: Amplifying Creative Ability… | HCOMP | 2014 | ||||||||||||||||||||||
| (LaToza et al., 2014) | Microtask programming: building… | UIST | 2014 | ||||||||||||||||||||||
| (Chilton et al., 2014) | Frenzy: Collaborative Data… | CHI | 2014 | ||||||||||||||||||||||
| (Weir et al., 2015) | Learnersourcing Subgoal Labels… | CSCW | 2015 | ||||||||||||||||||||||
| (Agapie et al., 2015) | Crowdsourcing in the Field: A… | HCOMP | 2015 | ||||||||||||||||||||||
| (Huang et al., 2015) | Guardian: A Crowd-Powered Spoken… | HCOMP | 2015 | ||||||||||||||||||||||
| (Cheng and Bernstein, 2015) | Flock: Hybrid Crowd-Machine… | CSCW | 2015 | ||||||||||||||||||||||
| (Lasecki et al., 2015) | Apparition: Crowdsourced User… | CHI | 2015 | ||||||||||||||||||||||
| (Luther et al., 2015) | Crowdlines: Supporting Synthesis… | HCOMP | 2015 | ||||||||||||||||||||||
| (Drapeau et al., 2016) | MicroTalk: Using Argumentation… | HCOMP | 2016 | ||||||||||||||||||||||
| (Hahn et al., 2016) | The Knowledge Accelerator: Big… | CHI | 2016 | ||||||||||||||||||||||
| (Kim and Monroy-Hernandez, 2016) | Storia: Summarizing Social Media… | CSCW | 2016 | ||||||||||||||||||||||
| (McDonnell et al., 2016) | Why Is That Relevant?… | HCOMP | 2016 | ||||||||||||||||||||||
| (Nebeling et al., 2016) | WearWrite: Crowd-Assisted Writing… | CHI | 2016 | ||||||||||||||||||||||
| (Goto et al., 2016) | Understanding Crowdsourcing… | HCOMP | 2016 | ||||||||||||||||||||||
| (Guo et al., 2016) | VizLens: A Robust and Interactive… | UIST | 2016 | ||||||||||||||||||||||
| (Teevan et al., 2016) | Supporting Collaborative Writing… | CHI | 2016 | ||||||||||||||||||||||
| (Butler et al., 2017) | More than a Feeling: The MiFace… | UIST | 2017 | ||||||||||||||||||||||
| (Chang et al., 2017) | Revolt: Collaborative Crowdsourcing… | CHI | 2017 | ||||||||||||||||||||||
| (Guo et al., 2017) | Facade: Auto-generating Tactile… | CHI | 2017 | ||||||||||||||||||||||
| (Huang et al., 2017) | Supporting ESL Writing by… | HCOMP | 2017 | ||||||||||||||||||||||
| (Lee et al., 2017) | SketchExpress: Remixing Animations… | UIST | 2017 | ||||||||||||||||||||||
| (Salisbury et al., 2017) | Toward Scalable Social Alt Text:… | HCOMP | 2017 | ||||||||||||||||||||||
| (Valentine et al., 2017) | Flash Organizations: Crowdsourcing… | CHI | 2017 | ||||||||||||||||||||||
| (Kaur et al., 2017) | CrowdMask: Using Crowds to… | HCOMP | 2017 | ||||||||||||||||||||||
| (Kim et al., 2017) | Mechanical Novel: Crowdsourcing… | CSCW | 2017 | ||||||||||||||||||||||
| (Salehi et al., 2017) | Communicating Context to the… | CSCW | 2017 | ||||||||||||||||||||||
| (Liu et al., 2018) | ConceptScape: Collaborative… | CHI | 2018 | ||||||||||||||||||||||
| (Huang et al., 2018) | Evorus: A Crowd-powered… | CHI | 2018 | ||||||||||||||||||||||
| (Gouravajhala et al., 2018) | EURECA: Enhanced Understanding… | HCOMP | 2018 | ||||||||||||||||||||||
| (Mahyar et al., 2018) | CommunityCrit: Inviting the… | CHI | 2018 | ||||||||||||||||||||||
| (Kobayashi et al., 2018) | An Empirical Study on Short-… | HCOMP | 2018 | ||||||||||||||||||||||
| (Guo et al., 2019) | StateLens: A Reverse Engineering… | UIST | 2019 | ||||||||||||||||||||||
| (Mohanty et al., 2019) | Second Opinion: Supporting… | HCOMP | 2019 | ||||||||||||||||||||||
| (Chilton et al., 2019) | VisiBlends: A Flexible Workflow… | CHI | 2019 | ||||||||||||||||||||||
| (Chen et al., 2019) | Cicero: Multi-Turn, Contextual… | CHI | 2019 | ||||||||||||||||||||||
| (Alshaibani et al., 2020) | Privacy-Preserving Face Redaction… | HCOMP | 2020 | ||||||||||||||||||||||
| (Huffaker et al., 2020) | Crowdsourced Detection of… | CHI | 2020 | ||||||||||||||||||||||
| (Bragg et al., 2021) | ASL Sea Batle: Gamifying Sign… | CHI | 2021 | ||||||||||||||||||||||
| (Alshaibani and Quinn, 2021) | Pterodactyl: Two-Step Redaction… | HCOMP | 2021 |
Reference |
Title |
Venue |
Year |
Quality Increase |
Quality Decrease |
Cost Increase |
Cost Decrease |
Objectivity |
Subjectivity |
Task General |
Task Specific |
Sequential |
Branching |
Redundant |
Dynamic |
Communicative |
Generate |
Evaluate |
Improve |
Focus |
Partition |
Merge |
Crowdworkers |
LLMs |
User |
| (Cobbe et al., 2021) | Training Verifiers to Solve Math… | arxiv | 2021 | ||||||||||||||||||||||
| (Wu et al., 2022a) | PromptChainer: Chaining Large… | CHI | 2022 | ||||||||||||||||||||||
| (Bursztyn et al., 2022) | Learning to Perform Complex… | EMNLP | 2022 | ||||||||||||||||||||||
| (Eisenstein et al., 2022) | Honest Students from Untrusted… | arxiv | 2022 | ||||||||||||||||||||||
| (Wu et al., 2021) | AI Chains: Transparent and… | CHI | 2022 | ||||||||||||||||||||||
| (Zhang et al., 2022) | AUTOMATIC CHAIN OF THOUGHT… | ICLR | 2022 | ||||||||||||||||||||||
| (Creswell and Shanahan, 2022) | Faithful Reasoning Using Large… | arxiv | 2022 | ||||||||||||||||||||||
| (Creswell et al., 2022) | Selection-Inference: Exploiting… | arxiv | 2022 | ||||||||||||||||||||||
| (Levine et al., 2022) | Standing on the Shoulders of… | arxiv | 2022 | ||||||||||||||||||||||
| (Schick et al., 2022) | PEER: A Collaborative Language… | arxiv | 2022 | ||||||||||||||||||||||
| (Wang et al., 2022) | Iteratively Prompt Pre-trained… | EMNLP | 2022 | ||||||||||||||||||||||
| (Park et al., 2022) | Social Simulacra: Creating… | UIST | 2022 | ||||||||||||||||||||||
| (Yang et al., 2022) | Re3: Generating Longer Stories… | EMNLP | 2022 | ||||||||||||||||||||||
| (Reppert et al., 2023) | ITERATED DECOMPOSITION: IMPROVING… | arxiv | 2023 | ||||||||||||||||||||||
| (Arora et al., 2023) | Ask Me Anything: A simple strategy… | ICLR | 2023 | ||||||||||||||||||||||
| (Huang et al., 2023c) | Ai chain on large language model… | arxiv | 2023 | ||||||||||||||||||||||
| (Parameswaran et al., 2023) | Revisiting Prompt Engineering… | arxiv | 2023 | ||||||||||||||||||||||
| (Wang et al., 2023) | PopBlends: Strategies for… | CHI | 2023 | ||||||||||||||||||||||
| (Xie et al., 2023) | Decomposition Enhances Reasoning… | arxiv | 2023 | ||||||||||||||||||||||
| (Yao et al., 2024) | Tree of Thoughts: Deliberate… | Neurips | 2023 | ||||||||||||||||||||||
| (Zharikova et al., 2023) | DeepPavlov Dream: Platform for… | ACL | 2023 | ||||||||||||||||||||||
| (Fleming et al., 2023) | Assessing the Potential of… | medRxiv | 2023 | ||||||||||||||||||||||
| (Huang et al., 2023a) | Adaptive Intellect Unleashed:… | CORR | 2023 | ||||||||||||||||||||||
| (Li et al., 2023) | Teach LLMs to Personalize – An… | arxiv | 2023 | ||||||||||||||||||||||
| (Long et al., 2023) | Tweetorial Hooks: Generative AI… | ICCC | 2023 | ||||||||||||||||||||||
| (Nair et al., 2023a) | Generating medically-accurate… | ACL | 2023 | ||||||||||||||||||||||
| (Wu et al., 2023a) | AutoGen: Enabling Next-Gen LLM… | arxiv | 2023 | ||||||||||||||||||||||
| (Zhou et al., 2022b) | LEAST-TO-MOST PROMPTING ENABLES… | ICLR | 2023 | ||||||||||||||||||||||
| (Huang et al., 2023b) | PCR-Chain: Partial Code Reuse… | ICSE | 2023 | ||||||||||||||||||||||
| (Gero et al., 2023) | Self-Verification Improves… | arxiv | 2023 | ||||||||||||||||||||||
| (Kim et al., 2023b) | Metaphorian: Leveraging Large… | DIS | 2023 | ||||||||||||||||||||||
| (Liang et al., 2023) | Encouraging Divergent Thinking… | arxiv | 2023 | ||||||||||||||||||||||
| (Madaan et al., 2024) | Self-Refine: Iterative Refinement… | Neurips | 2023 | ||||||||||||||||||||||
| (Mirowski et al., 2023) | Co-Writing Screenplays and… | CHI | 2023 | ||||||||||||||||||||||
| (Zhang et al., 2023a) | VISAR: A Human-AI Argumentative… | UIST | 2023 | ||||||||||||||||||||||
| (Du et al., 2023) | Improving Factuality and Reasoning… | arxiv | 2023 | ||||||||||||||||||||||
| (Nair et al., 2023b) | DERA: Enhancing Large Language… | arxiv | 2023 | ||||||||||||||||||||||
| (Wu et al., 2023b) | LLMs as Workers in Human-Computational… | arxiv | 2023 | ||||||||||||||||||||||
| (Wu et al., 2024) | An Empirical Study on Challenging… | ICLR Workshop | 2024 |
4.3. Comparing the utilization of techniques
This section discusses the distribution of the crowdsourcing workflow and LLM chaining literature across the design space. We fit each paper in the corpus into its design space categories. The labeling procedure and outcomes are detailed in Supplementary materials and in Tables 1 and 2. Here we provide a snapshot of the LLM chaining sub-field as it expands, but these distributions will change over time with more published papers.
Comparing these distributions first shows how both literatures cover all design space categories, with the natural exception of the LLM and crowdworker categories. The methods for creating the design space did not enforce overlap of the literatures, so the high degree of overlap between them provides a sanity check that both fields can share a design space.
However, their distribution differs across some categories. For instance, a greater proportion of crowdsourcing papers promote the decrease of cost compared to experts, while LLM chains’ costs are usually compared to a baseline of one call to the LLM, inherently increasing the costs. Due to the high range of tasks and sparsity of reporting, we cannot compare the scale of the relative costs of workflows and chains. Furthermore, differences in execution details hinder comparison across the literatures even for similar tasks.
Across the objectivity-subjectivity spectrum, we find that crowdsourcing workflows more often implement a task with subjectivity, while LLM chains more often have purely objective tasks such as math. Finally, more chains in the LLM literature are designed to be generalizable to many tasks, while a greater proportion of workflows in the crowdsourcing literature are designed for a specific task function.
The crowdsourcing workflow literature uses redundant and especially communicative architectures more frequently. More chaining papers use only a sequential architecture than crowdsourcing papers (9 papers vs. 1 paper). There are slight differences in subtask use, in which the crowdsourcing workflow literature uses evaluate and improve subtasks more often, and the LLM chaining literature uses generate, focus and partition subtasks more often. Finally, more papers in the crowdsourcing workflow literature explicitly incorporate the user than those in the LLM chaining literature.
Three columns headed by ”If…”, ”Then…” and ”With…”. The ”If” column contains different possible challenges relating to achieving objectives. The ”Then” column lists all of the strategies. There are arrows connecting the ”If” column to the ”Then” column, with arrows from ”navigating a large creative space” objectives connect to the adaptive architectures, diverse responses, and user guidance strategies. The ”With” column lists all of the tactics. There are arrows connecting the ”Then” column to the ”With” column. There are many arrows, and they correspond to the tactics used for strategies as defined in the text and other figures.
Surfacing the differences between these distributions focuses attention on potentially informative sections of the design space. These differences could result for a multitude of reasons. We speculate that the lower rate of subjective tasks in chaining reflects a relative preference for benchmark evaluations in the LLM chaining community. The incorporation of the user and the focus on task specificity could reflect that crowdsourcing work was primarily housed in the Human-Computer Interaction field. The variations in these distributions could also reflect inherent differences between LLMs and crowdworkers or indicate areas of potential future work.
Comparing the two distributions confirms the many similarities between crowdsourcing and chaining methods while focusing attention on the explanatory and actionable potential of the differences. We explore these differences in further detail in the case studies (Section 7) and the discussion (Section 8).
@@ 5. Strategies integrating the design space
Choosing tactic designs to support objectives is a non-trivial task, as combining different tactic types and implementations supports a huge range of possible workflows. However, through the thematic coding process we surfaced strategies by which tactics are used to support objectives (Figure 5). Specifically, these strategies (5.1) encourage diverse responses, (5.2) validate outputs from subtasks, (5.3) enable user guidance, (5.4) utilize adaptable architectures, (5.5) improve actor selection, (5.6) align a subtask’s required skills with an actor’s capabilities, (5.7) promote global context, and (5.8) ensure outputs meet quality thresholds.
@@
Title. Diverse responses, to promote objective or subjective quality, source multiple responses to the same subtask with parallel, iterative, or debate based workflows. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Robust to low quality responses, Coherency and cascading errors), Constraints (increase expense and latency), Objectivity (Robust to inaccurate outputs), Subjectivity (Improve with divergent thinking). Section 2: tactics used. Actors (LLMs, crowdworkers), Subtasks (all), Architectures (redundant, communicative).
@@ 5.1. Diverse responses
To promote objective or subjective quality, source multiple responses to the same subtask with parallel, iterative, or communicative workflows (Figure 6).
@@
Motivation. Individual actors can give low quality outputs, affecting the quality of the overall output. Designers may want to improve robustness to low quality outputs, outliers, and spammers to support the accuracy of objective outcomes [C(Liem et al., 2011; Lasecki et al., 2013; Huang et al., 2018; Hahn et al., 2016; Zaidan and Callison-Burch, 2011; Ambati et al., 2012; Huffaker et al., 2020) L(Du et al., 2023; Reppert et al., 2023)]. For subjective tasks, they may want to encourage divergent and diverse thinking [C(Luther et al., 2015; Lasecki et al., 2015; Nebeling et al., 2016; Huang et al., 2017; Teevan et al., 2016) L(Wu et al., 2023a; Liang et al., 2023)].
@@
Implementation. Getting responses from multiple actors on a subtask can help support these goals. Crowdsourcing workflows and LLM chaining source diverse responses differently. Crowdsourcing does so by querying different people. Chaining leverages the stochastic nature of LLMs (e.g., with a higher temperature parameter setting) or uses different models or prompt variants (e.g., role-based prompts) to source a variety of responses [L(Wu et al., 2023b; Park et al., 2022; Arora et al., 2023)].
@@
Redundant architectures are the most direct way of sourcing diverse responses, but communicative architectures, such as those that structure debates, can source diverse responses as well. Some studies in crowdsourcing systematically study the impact of different architectures (e.g., parallel vs iterative) and architecture parameters (e.g., the number of iterations) on the diversity and quality of the outcomes [C(André et al., 2014; Kittur et al., 2011; Little et al., 2010a; Goto et al., 2016)]. However, it is unknown if these findings transfer to similar tactic choices with LLMs.
@@
Secondary benefits. There are some benefits of diverse responses that are solely discussed in the crowdsourcing literature, as they are dependent on the variations among people. Crowdsourcing elicits the benefit of having different viewpoints [C(Lasecki et al., 2013; Ambati et al., 2012; Chen et al., 2019)], filling in other crowdworkers’ deficits to create a more rounded final product [C(Zhang et al., 2012; Kim and Monroy-Hernandez, 2016)], and sourcing perspectives from expert crowdworkers [C(Chilton et al., 2014)] or from crowdworkers with various skill levels [C(Mahyar et al., 2018)].
@@
Consequences. Sourcing diverse responses is inherently more costly because multiple actors complete the same subtask. Merging multiple responses can induce outcome quality errors: coherency for parallel architectures [L(Wu et al., 2022a)] and cascading errors for iterative architectures [C(Goto et al., 2016)].
@@
Examples. Examples of the diverse responses strategy include: Apparition [C(Lasecki et al., 2015)], chains for multi-agent debate [L(Liang et al., 2023)], and an exploration of ”human computation processes” [C(Little et al., 2010a)].